OpenStudy (denonakavro):

AP Stats help !!! will medal and fan The University of Washington claims that it graduates 85% of its basketball players. An NCAA investigation about the graduation rate finds that of the last 20 players entering the program only 11 graduated. If the university’s claim is correct, the number of players among the 20 should follow the binomial distribution B(20, 0.85). a) What is the probability that 11 out of the 20 would graduate? Write how the binomial formula is used to calculate the probability. I did it with my calculator and showed it to my teacher, he said I've to use bin formula

OpenStudy (anonymous):

Well, the binomial probability for \(r\) successes in \(n\) trials (each with probability \(p\)) is \[B(r,n,p) = {n\choose r} \cdot p^r (1-p)^{n-r} \] where \[{n\choose r} = \frac{n!}{r!(n-r)!} \]